OpenStudy (anonymous):

At the start of a basketball game, the referee tosses a ball into the air for the jump ball. The equation that models the height of the ball is h(t) = -16t2 + 30t + 8. During what time interval is the height of the ball above 10 feet? (Round your answer to the nearest tenth of a second.) a) Between 1.1 and 1.6 seconds b) Between 0.1 and 1.8 seconds c) Between 0.3 and 2.1 seconds

OpenStudy (anonymous):

Well, \(-16t^{2}+30t+8=10 => -16t^{2} +30t-2=0\) Then solve for the time \(t\)

OpenStudy (anonymous):

Using the quadratic formula.

OpenStudy (anonymous):

\[\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2a} \]

OpenStudy (anonymous):

I guest have no choice but, to learn how to master Quadratic Formula

OpenStudy (anonymous):

You don't know how to use the Quadratic FOrmula? It's basically subbing in numbers.

OpenStudy (anonymous):

b)Between 0.1 and 1.8 seconds. is my educate guess i lost trace